# Neural Transparency: Mechanistic Interpretability Interfaces for Anticipating Model Behaviors for Personalized AI

### v1

@@ 3.1.2. Extracting Personality from Responses of Target LLM
Following an evaluation of several open-source large language models, we selected Llama-3.2-3B-Instruct based on its compact architecture, which enabled responsive real-time interaction and reduced the computational overhead associated with persona vector generation while still exhibiting rich and nuanced conversational capabilities. All combinations of system prompts and extraction question prompts were passed into Llama to create 400 unique responses, and the activations from each forward pass were cached using Transformer Lens (Nanda and Bloom, 2022).
@@
The resulting activations for each kept response were of shape (num_layers, num_tokens, hidden_dim). We computed the mean across all tokens within each response, reducing the token dimension to yield activations of shape (num_layers, hidden_dim). Next, we averaged these activations across all kept responses separately for the positive and negative system prompt conditions, producing two mean activation tensors. The persona vector was then computed as the difference between these contrastive representations. By subtracting the mean negative activations from the mean positive activations, we obtained a vector that captures the linear direction in activation space along which the target trait is represented.
@@ 3.2.1. Using Projection to Calculate Persona Scores
We used the persona vectors to construct persona scores, which quantify the predicted level of trait expression for an LLM chatbot given a custom system prompt. For a given trait, the persona score $s$ of a system prompt was computed by projecting the activation vector of the final token of the system prompt $\mathbf{a}\in\mathbb{R}^{d}$ onto the corresponding persona vector $\mathbf{b}\in\mathbb{R}^{d}$
@@
3.2.2. Optimizing Persona Scores through Neural Layer Selection
@@
For each trait, we generated five system prompts at each of the five trait expression levels, yielding 25 synthetic prompts per trait. We then computed persona scores using activations from each layer and performed linear regression between the specified trait expression level and the corresponding persona score. The layer yielding the highest mean $R^{2}$ value across all traits was selected for subsequent analyses, as this was the layer that best predicted trait expression and measured a linear representation. We identified layer 20 of 26 in Llama-3.2-3B-Instruct as the best layer. Figure 3 visualizes the values of the activations in layer 20. The activations of the layer were shaped into a two-dimensional representation for visualization purposes. This regression analysis additionally quantified the predictive validity of our persona scores for trait expression.
@@ 3.3. Persona Vector Evaluation
Our analysis revealed linear relationships between trait expression levels in LLM-generated system prompts and their corresponding persona scores, validating our persona vector generation method. The hallucination trait exhibited a notably weak relationship, which we attribute to its behavioral complexity. Unlike simple emotive dimensions that exist on clear semantic binaries (empathy, sociality, formality), hallucination is a behavior that can be difficult to consistently elicit or occur predictably.
This behavioral complexity likely produced more variance in the LLM responses in the persona generation pipeline, yielding less representative activation samples. Additionally, safety mechanisms in the LLMs may have led to less accurate representations for safety-relevant traits. Despite filtering out responses where the model refused to answer, higher refusal rates for these traits would have reduced available training examples and potentially reduced the faithfulness of the resulting persona vectors.
@@ 3.4. Neural Transparency Interface Design
The sunburst visualization enables us to present the persona scores in an intuitive and visually appealing manner. It is designed in a radial layout containing two concentric rings that encodes both categorical and quantitative information about the chatbot’s personality traits resulting from the user’s system prompt (Figure 5). The user also chooses an avatar that is placed in the center of the sunburst to represent their custom AI personality. We developed the visualization using the JavaScript visualization library, D3 (Bostock et al., 2011), and the web interface using a mixture of JavaScript, HTML, and CSS.
@@ 3.6. Participants and Ethics
We recruited eighty participants using the online hosting platform Prolific to engage in a 30-minute study where they were asked to create an AI companion that can provide emotional support in difficult times. We selected participants that were from the US, English-speaking, and owned a laptop or desktop. The ages of the participants ranged from 20 to 69 years (M=42.3, SD=11.4), with 44 participants identifying as male and 36 identifying as female. The protocol was reviewed and granted an exemption by a [Redacted] Institutional Review Board.
@@ 3.9.2. Effect of Neural Transparency on Design Iteration, AI Behavior, and User Engagement
AI Behavior Changes. To assess whether the visualization condition influenced the types of personalities participants created, we analyzed trait-level changes in the AI companions. Specifically, we: (1) calculated persona scores for both the initial and final system prompts, (2) computed the difference between these scores for each personality trait, (3) compared these trait shifts between conditions to identify whether visualization led participants to activate different personality characteristics.
@@ 4. Results
We find evidence suggesting that users systematically miscalibrate how their personalized AI will behave, consistently over-estimating or under-estimating trait expressions across most dimensions (eleven of fifteen analyzable traits, all p ¡ .05). This fundamental miscalibration demonstrates that users may not reliably anticipate model behavior from system prompts alone, warranting the development of mechanistic interpretability interfaces.
@@
Nonetheless, neural transparency significantly impacted user trust and perception. Users who received neural transparency feedback reported significantly higher trust in their AI companion (p = .042, Cohen’s d = 0.46) and rated the visualization as highly helpful (M = 5.98/7). Remarkably, most participants expressed strong desire to use such tools again in future chatbot design (M = 6.05/7). These findings suggest that while our current interface design did not translate transparency into measurable behavioral improvements, users found value in understanding their AI’s internal representations, a disconnect that points to important directions for refining the interfaces in the future.
@@ 4.1. Behavioral anticipation accuracy
Participants were asked to predict trait expressions from system prompts across eight binary trait dimensions (16 total trait expressions). We conducted independent samples t-tests to understand potential baseline differences between the two conditions on these 0–10 ratings of predicted trait expression. We found that there were no group differences in how participants rated any of the trait dimensions ($p>.05$ for all eight dimensions). This establishes that each group had no prior bias in their expectations—both groups believed their system prompts would elicit similar emotional-support behaviors from the chatbot.
@@ 4.1.1. Human mental models of trait activations
We sought to understand if participants could extrapolate accurate representations of the persona vectors from their system prompts alone. Despite participants’ predictions being positively correlated with actual trait activations across all analyzable traits ($p<.001$ for all fifteen traits), participants consistently misestimated the degree to which traits would be activated from their system prompts.
@@
Conversely, participants significantly underestimated the activation of one negatively valenced trait, sycophantic (t(79)=3.789t(79)=-3.789, $p<.001$, d=0.424d=-0.424). Similarly, some traits of neutral valence such as formal (t(79)=3.147t(79)=-3.147, $p=.002$, d=0.352d=-0.352) and serious (t(79)=16.286t(79)=-16.286, $p<.001$, d=1.821d=-1.821), were significantly underestimated. The underestimation of “serious” was particularly pronounced, representing the largest effect size in the analysis (d=1.821d=-1.821).
Two traits showed marginal or small but significant effects: respectful (t(79)=2.615t(79)=-2.615, $p=.011$, d=0.169d=-0.169) and antisocial ($t(79)=2.093$, $p=.040$, $d=0.234$), indicating that participants were marginally inaccurate in predicting the consequences of their system prompts on these traits. Only four traits showed no significant difference between predicted and actual activations: unempathetic ($t(79)=0.095$, $p=.924$, $d=0.011$), social ($t(79)=0.240$, $p=.811$, $d=0.027$), hallucinatory ($t(79)=1.569$, $p=.121$, $d=0.175$), and discouraging (t(79)=1.512t(79)=-1.512, $p=.135$, d=0.169d=-0.169), suggesting accurate calibration for these specific dimensions.
@@ 4.2. Effect of Visualization on User Engagement and Design Iteration
Participants also showed similar patterns of design iteration during the prompt refinement phase. The number of unique prompts generated did not significantly differ between control ($M=1.58$, $SD=1.00$) and experimental conditions ($M=1.64$, $SD=1.06$), t(78)=0.277t(78)=-0.277, $p=.783$, Cohen’s d=0.06d=-0.06. Similarly, the number of persona scores generated was comparable across control ($M=1.61$, $SD=1.00$) and experimental conditions ($M=1.79$, $SD=1.12$), t(78)=0.758t(78)=-0.758, $p=.451$, Cohen’s d=0.17d=-0.17. These negligible effect sizes suggest that the experimental manipulation did not meaningfully influence how extensively participants iterated on their AI designs [addressed in more detail in Discussion].
4.3. Effect of Visualization on AI Behaviors
We examined whether the visualization affected the types of personas participants created by calculating the difference between the first and last persona scores (delta scores). Independent samples t-tests revealed that personality traits did not shift significantly differently between control and experimental conditions across any of the measured dimensions.
We found no significant differences in changes toward empathetic, unempathetic, encouraging, discouraging, casual, formal, serious, factual, hallucinatory, sycophantic, anti-social, or respectful ($p>0.05$). These findings suggest that the visualization did not systematically influence the direction or magnitude of personality changes participants made to their AI chatbots during the design process.
@@ 4.4.1. Baseline Equivalence and Post-Interaction Assessments
Independent samples t-tests confirmed no significant baseline differences between conditions, allowing us to attribute differences post-interaction to the experimental manipulation. Prior to interacting with their chatbot, participants in the control and visualization conditions reported equivalent levels of trust (‘pre_trust’ in Figure 8), predictions about general unintended behaviors (‘pre_predict_unintended_behaviors’) and negative unintended behaviors (‘pre_predict_negative_behaviors’).
@@
We found that the neural transparency visualization significantly increased user trust in their chatbot. Participants in the visualization condition reported higher trust ($M=5.60$, $SD=0.91$) compared to the control condition ($M=5.13$, $SD=1.10$), representing a small-to-medium effect (t(78)=2.065t(78)=-2.065, $p=.042$, Cohen’s $d=0.46$). Notably, while post-interaction trust differed between groups, the change in the magnitude of trust from pre-interaction to post-interaction did not differ (‘delta_trust’ in Figure 8), suggesting that both groups experienced similar trajectories of trust development but arrived at different endpoints.
@@
After using the chat, participants in both conditions reported similar confidence in their ability to predict unintended behaviors. No significant differences emerged for predicting general unintended behaviors (‘post_predict_unintended_behaviors’ in Figure 8) or specifically negative unintended behaviors (‘post_predict_negative_behaviors’). Similarly, changes in prediction confidence from pre-interaction to post-interaction showed no significant differences between conditions for either general or negative unintended behaviors (‘delta_unintended’, ‘delta_neg_unintended’, respectively).
@@
Independent of condition, participants reported high satisfaction with their chatbot design. Both control ($M=5.97$, $SD=1.00$) and visualization ($M=6.21$, $SD=1.12$) groups felt they successfully arrived at their desired chatbot character, with no significant difference between conditions (‘post_arrived_desired_character’). This indicates that the neural transparency visualization did not make it more difficult for users to achieve a specific persona design or exceptionally enrich the design process.
@@ 4.4.2. Perception of the Visualization
When asked whether they would want to see the visualization again in future chatbot design tasks, participants responded very enthusiastically ($M=6.05$, $SD=1.32$), suggesting strong user acceptance and a clear desire to continue using mechanistic interpretability tools in AI companion creation. The high mean (approaching the maximum of 7) indicates that participants found lasting value in the transparency mechanism beyond novelty.
@@ 4.5.1. Understanding the Consequences of System Prompts
A key challenge for participants in the control condition was uncertainty about whether their system prompts were actually influencing the chatbot’s behavior. One control participant captured this frustration:
“It seemed like my prompt wasn’t followed nearly as closely as I would’ve liked and after a few changes, I kind of just left the prompt in the final stage to have a conversation and test it out.”
This ambiguity led some participants to abandon iterative refinement of their prompts. Another control participant noted similar difficulties despite multiple attempts:
“After I modified my prompt to include giving concise responses, the AI agent was concise at first, but then went back to giving long responses, which was frustrating.”
@@ 4.5.2. Users Needed More Interaction Time
An experimental participant similarly noted the abruptness of the time constraint:
“Everything worked exactly like I expected it. It was over a lot faster than I expected and would have loved a 2 minutes left to interact timer or something instead of just cutting off out of nowhere.”
@@ 5. Discussion
5.1. Inaccurate Mental Models: Why Users Need Transparency Tools
@@
Our most striking finding is that participants consistently incorrectly predicted how their system prompts would manifest in the personality assessment from the persona scores. Despite correlations between predictions and actual trait persona scores, participants systematically over-estimated or under-estimated trait expression for eleven of the fifteen analyzable trait expressions. This miscalibration occurred even though participants were designing chatbots for a specific purpose (emotional support) and had explicit intentions about desired behaviors.
@@
These inaccurate mental models provide strong motivation for transparency mechanisms in chatbot creation interfaces. Users cannot reliably anticipate emergent behaviors from system prompts alone, even when they understand their goals clearly. The opaqueness of this mechanism means users are essentially operating blind during the design process, discovering problems only after deployment through trial and error. This discovery-oriented approach is inefficient and potentially dangerous when chatbots are being created for vulnerable users or sensitive contexts.
The consistency of incorrect prediction across multiple traits suggests this is not simply a matter of users lacking domain knowledge about specific behavioral dimensions. Even users who successfully created satisfactory chatbots (as evidenced by high design satisfaction scores) demonstrated poor predictive accuracy, suggesting that achieving desired outcomes through iteration does not necessarily improve their mental models of the underlying system.
5.2. The Transparency Paradox: High Value to Users, Null Behavioral Effects

### v2

@@ 3.1.2. Extracting Personality from Responses of Target LLM
Following an evaluation of several open-source LLMs, we selected Llama-3.2-3B-Instruct based on its compact architecture, which enabled responsive real-time interaction and reduced the computational overhead associated with persona vector generation while still exhibiting rich and nuanced conversational capabilities. All combinations of system prompts and extraction question prompts were passed into Llama to create 400 unique responses, and the activations from each forward pass were cached using Transformer Lens (nanda2022transformerlens).
@@
The resulting activations for each kept response were of shape (num_layers, num_tokens, hidden_dim). We computed the mean across all tokens within each response, collapsing the token dimension to yield activations of shape (num_layers, hidden_dim). Next, we averaged these activations across all kept responses separately for the positive and negative system prompt conditions, producing two mean activation tensors. The persona vector was then computed as the difference between these contrastive representations. By subtracting the mean negative activations from the mean positive activations, we obtained a vector that captures the linear direction in activation space along which the target trait is represented.
@@ 3.2.1. Using Projection to Calculate Persona Scores
We used the persona vectors to construct persona scores, which quantify the predicted level of trait expression for an LLM chatbot given a custom system prompt. For a given trait, the persona score $s$ of a system prompt was computed by projecting the activations of the final token of the system prompt $\mathbf{a}\in\mathbb{R}^{d}$ onto the corresponding persona vector $\mathbf{b}\in\mathbb{R}^{d}$
@@
3.2.2. Optimizing Persona Scores through Layer Selection
@@
For each trait, we generated five system prompts at each of the five trait expression levels, yielding 25 synthetic prompts per trait. We then computed persona scores using activations from each layer and performed linear regression between the specified trait expression level and the corresponding persona score. The layer yielding the highest mean $R^{2}$ value across all traits was selected for subsequent analyses, as this was the layer that best predicted trait expression and measured a linear representation. We identified layer 20 of 26 in Llama-3.2-3B-Instruct as the best layer. Figure 3 visualizes the values of the activations in layer 20. The activations of the layer were shaped into a two-dimensional representation in our visualizations. This regression analysis additionally quantified the predictive validity of our persona scores for trait expression.
@@ 3.3. Persona Vector Evaluation
Our analysis revealed linear relationships between trait expression levels in LLM-generated system prompts and their corresponding persona scores, validating our persona vector generation method. The hallucination trait exhibited a notably weak relationship, which we attribute to its behavioral complexity. Unlike simple emotive dimensions that exist on clear semantic binaries (empathy, sociality, formality), hallucination is a behavior that can be difficult to consistently elicit or occur predictably. This behavioral complexity likely produced more variance in the LLM responses in the persona generation pipeline, yielding less representative activation samples. Additionally, safety mechanisms in the LLMs may have led to less accurate representations for safety-relevant traits. Despite filtering out responses where the model refused to answer, higher refusal rates for these traits would have reduced available training examples and potentially reduced the faithfulness of the resulting persona vectors.
@@ 3.4. Neural Transparency Interface Design
The sunburst visualization enables us to present the persona scores in an intuitive and visually appealing manner. It is designed in a radial layout containing two concentric rings that encodes both categorical and quantitative information about the chatbot’s personality traits resulting from the user’s system prompt (Figure 5). The user also chooses an avatar that is placed in the center of the sunburst to represent their custom AI personality. We developed the visualization using the JavaScript visualization library, D3 (bostock2011d3), and the web interface using a mixture of JavaScript, HTML, and CSS.
@@ 3.6. Participants and Ethics
We recruited eighty participants using the online hosting platform Prolific to engage in a 30-minute study where they were asked to create an AI companion that can provide emotional support in difficult times. We selected participants that were from the US, English-speaking, and owned a laptop or desktop. The ages of the participants ranged from 20 to 69 years ($M=42.3$, $SD=11.4$), with 44 participants identifying as male and 36 identifying as female. The protocol was reviewed and granted an exemption (E-7192) by the Massachusetts Institute of Technology Committee on the Use of Humans as Experimental Subjects (COUHES).
@@ 3.9.2. Effect of Neural Transparency on Design Iteration, AI Behavior, and User Engagement
AI Behavior Changes. To assess whether the visualization condition influenced the types of personalities participants created, we analyzed trait-level changes in the AI companions. Specifically, we calculated the final persona scores from system prompts and compared these scores between conditions.
@@ 4. Results
We find evidence suggesting that users systematically miscalibrate how their personalized AI will behave, consistently over-estimating or under-estimating trait expressions across most dimensions (eleven of fifteen analyzable traits, all p ¡ .05). This miscalibration demonstrates that users may not reliably anticipate model behavior from system prompts alone, warranting the development of mechanistic interpretability interfaces.
@@
Nonetheless, neural transparency significantly impacted user trust and perception. Users who received neural transparency feedback reported significantly higher trust in their AI companion (p = .042, Cohen’s d = 0.46) and rated the visualization as highly helpful (M = 5.98/7). Remarkably, most participants expressed strong desire to use such tools again in future chatbot design (M = 6.05/7). These findings suggest that while our current interface design did not translate transparency into measurable behavioral improvements, users found value in understanding their AI’s internal representations, a divergence that points to important directions for refining the interfaces in the future.
@@ 4.1. Behavioral Anticipation Accuracy
Participants were asked to predict the level of trait expressions from their system prompts across eight binary trait dimensions (16 total trait expressions). We conducted independent samples t-tests to understand potential baseline differences between the two conditions on ratings of 0–10 of predicted trait expression. We found that there were no group differences in how participants rated any of the trait dimensions ($p>.05$ for all eight dimensions). This establishes that each group had no prior differences in their expectations—both groups believed their system prompts would elicit similar emotional-support behaviors from the chatbot.
@@ 4.1.1. Human mental models of trait activations
We sought to understand if participants could accurately predict the persona scores from their system prompts alone. Despite participants’ predictions being positively correlated with actual trait activations across all analyzable traits ($p<.001$ for all fifteen traits), participants consistently misestimated the degree to which traits would be activated from their system prompts.
@@
Conversely, participants significantly underestimated the activation of one negatively valanced trait, sycophantic (t(79)=3.789t(79)=-3.789, $p<.001$, d=0.424d=-0.424). Similarly, some traits of neutral valence such as formal (t(79)=3.147t(79)=-3.147, $p=.002$, d=0.352d=-0.352) and serious (t(79)=16.286t(79)=-16.286, $p<.001$, d=1.821d=-1.821), were significantly underestimated. The underestimation of “serious” was particularly pronounced, representing the largest effect size in the analysis (d=1.821d=-1.821).
Two traits showed marginal or small but significant effects: respectful (t(79)=2.615t(79)=-2.615, $p=.011$, d=0.169d=-0.169) and antisocial ($t(79)=2.093$, $p=.040$, $d=0.234$), indicating that participants had small prediction errors regarding the consequences of their system prompts on these persona traits. Only four traits showed no significant difference between predicted and actual activations: unempathetic ($t(79)=0.095$, $p=.924$, $d=0.011$), social ($t(79)=0.240$, $p=.811$, $d=0.027$), hallucinatory ($t(79)=1.569$, $p=.121$, $d=0.175$), and discouraging (t(79)=1.512t(79)=-1.512, $p=.135$, d=0.169d=-0.169), suggesting accurate calibration for these specific dimensions.
@@ 4.2. Effect of Visualization on User Engagement and Design Iteration
Participants also showed similar patterns of design iteration during the prompt refinement phase. The number of unique prompts generated did not significantly differ between control ($M=1.58$, $SD=1.00$) and experimental conditions ($M=1.64$, $SD=1.06$), t(78)=0.277t(78)=-0.277, $p=.783$, Cohen’s d=0.06d=-0.06. Similarly, the number of persona scores generated was comparable across control ($M=1.61$, $SD=1.00$) and experimental conditions ($M=1.79$, $SD=1.12$), t(78)=0.758t(78)=-0.758, $p=.451$, Cohen’s d=0.17d=-0.17. These negligible effect sizes suggest that the experimental manipulation did not meaningfully influence how extensively participants iterated on their AI designs.
4.3. Effect of Visualization on AI Personalities
We utilized independent-samples t-tests to examine whether the visualization affected the users’ final chatbot personalities by comparing the persona scores for their final system prompts between experimental conditions. We found that the final chatbots had similar levels of expression for each trait ($p>0.05$), which indicates that participants ended up designing similar chatbot personalities between groups. It may also suggest that the visualization did not encourage users to selectively increase or decrease the expression of different traits with system prompting.
@@ 4.4.1. Baseline Equivalence and Post-Interaction Assessments
Independent samples t-tests confirmed no significant baseline differences between conditions at the beginning of the study, allowing us to attribute differences post-interaction to experimental effects. Prior to interacting with their chatbot, participants in the control and visualization conditions reported equivalent levels of trust (pre_trust in Figure 8), predictions about general unintended behaviors (pre_predict_unintended_behaviors) and negative unintended behaviors (pre_predict_negative_behaviors).
@@
We found that the neural transparency visualization significantly increased user trust in their chatbot. Participants in the visualization condition reported higher trust ($M=5.60$, $SD=0.91$) compared to the control condition ($M=5.13$, $SD=1.10$), representing a small-to-medium effect (t(78)=2.065t(78)=-2.065, $p=.042$, Cohen’s $d=0.46$). Notably, while post-interaction trust differed between groups, the change in the magnitude of trust from pre-interaction to post-interaction did not differ (delta_trust in Figure 8), suggesting that both groups experienced similar trajectories of trust development but arrived at different endpoints.
@@
After using the chat, participants in both conditions reported similar confidence in their ability to predict unintended behaviors. No significant differences emerged for predicting general unintended behaviors (post_predict_unintended_behaviors in Figure 8) or specifically negative unintended behaviors (post_predict_negative_behaviors). Similarly, changes in prediction confidence from pre-interaction to post-interaction showed no significant differences between conditions for either general or negative unintended behaviors (delta_unintended and delta_neg_unintended).
@@
Independent of condition, participants reported high satisfaction with their chatbot design. Both control ($M=5.97$, $SD=1.00$) and visualization ($M=6.21$, $SD=1.12$) groups felt they successfully arrived at their desired chatbot character, with no significant difference between conditions (post_arrived_desired_character). This indicates that the neural transparency visualization did not make it more difficult for users to achieve a specific persona design or exceptionally enrich the design process.
@@ 4.4.2. Perception of the Visualization
When asked whether they would want to see the visualization again in future chatbot design tasks, participants responded very enthusiastically ($M=6.05$, $SD=1.32$), suggesting strong user acceptance and a clear desire to continue using mechanistic interpretability tools in AI companion creation. The high mean score of 6.05 (near the maximum of 7) indicates that participants found lasting value in the transparency mechanism beyond novelty.
@@ 4.5.1. Understanding the Consequences of System Prompts
A key challenge for one participant in the control condition was uncertainty about whether their system prompts were actually influencing the chatbot’s behavior:
“It seemed like my prompt wasn’t followed nearly as closely as I would’ve liked and after a few changes, I kind of just left the prompt in the final stage to have a conversation and test it out. Overall, it wasn’t bad, but not ready for primetime.“
@@ 5. Discussion
5.1. Why Users Need Transparency Tools
@@
Our most interesting finding from the user study is that participants had consistent prediction errors for how their system prompts would manifest as persona scores. Despite correlations between predictions and actual trait persona scores, participants incorrectly estimated trait expression for eleven of the fifteen analyzable trait expressions. This miscalibration occurred even though participants were designing chatbots for a specific purpose (emotional support) and had explicit intentions about desired behaviors.
@@
These inaccurate mental models provide strong motivation for transparency mechanisms in chatbot creation interfaces. Users cannot reliably anticipate emergent behaviors from system prompts alone. The opaqueness of this mechanism means users are essentially operating blind during the design process, discovering problems only after deployment through trial and error. This discovery-oriented approach is inefficient and potentially dangerous when chatbots are being created for vulnerable users or sensitive contexts.
5.2. High Value to Users, Null Behavioral Effects